OpenStudy (adalrymple):

Which of the following shows the real number expressions in order from least to greatest? Select one: A. six minus square root of eight then three squared minus two then square root of sixteen plus zero point eight. B. three squared minus two then six minus square root of eight then square root of sixteen plus zero point eight. C. six minus square root of eight then square root of sixteen plus zero point eight then three squared minus two. D. square root of sixteen plus zero point eight then six minus square root of eight then three squared minus two.

OpenStudy (legomyego180):

C

OpenStudy (adalrymple):

can u do the problem out please, I don't understand this @legomyego180

OpenStudy (legomyego180):

Sure, first off are you allowed to use calculators?

OpenStudy (adalrymple):

yes

OpenStudy (legomyego180):

Im going to reword the question. Put the following in order from smallest to largest: \[(6-\sqrt{8})...... (3^2-2)...... (\sqrt{16}+.8)\]

OpenStudy (legomyego180):

So all you need to do now is rewrite these as numbers you recognize then put them in order.

OpenStudy (legomyego180):

So type \[6-\sqrt{8}\] into your calculator and you get 3.17 Next \[3^2-2\] Thats just equal to 9-2 which is seven. So so far our numbers are 3.17, and 7 Last one \[\sqrt{16} + .8\] Well we know the square root of 16 is 4 and 4+.8 is 4.8 So our numbers are 3.17, 7 and 4.8 Now just put the numbers in order and list them out as they originally looked so it will be \[6-\sqrt{8}, \sqrt{16}+.8, 3^2-2\]